Calculate the break-even value using the marginal income ratio.

1.2 The sales manager proposes a R2 per unit reduction in selling price with the

expectation that this would increase sales by 2 000 units. Is this a good idea?

Motivate your answer.

(4 marks)

1.3 Determine the selling price per unit if a net profit of R486 000 is desired.


INFORMATION


The following forecasts for January 2021 were provided by Walton's Manufacturers:


Sales (36 000 units) ----------------------------------------R1080 000

Direct materials cost per unit ----------------------------R8

Direct labour cost per unit ---------------------------------R5

Variable manufacturing overhead costs per unit ----------R2

Fixed manufacturing overheads -------------------------------R150 000

Fixed marketing and administrative costs--------------------R120 000


Sales commission ------------------------------R10% OF SALES


1
Expert's answer
2020-10-05T18:31:10-0400

From the information given we need to do our calculation first in windows excel as shown below





"Q_1" Calculate the break even value using the marginal income ratio?


"BEP=\frac{total fixed cost}{contribution margin ratio}"


="\frac{270000}{40}=R675000"


"Q_2" No, this would decrease the income by R44,400.This is because the total sales went down while the expenses remained almost the same , making an overall reduction of profit


sales went down from 1,080,000 to 1,064,000 while expenses were .The profit generated in the first case was

"162,000" while in the second case was 117,600. So there was a reduction"(162000-117600)=44400" of this amount on profit


"Q_3" selling price per unit if a net profit of R486000 is desired


="=\frac{total sales}{number of units}=\frac{1344600}{36000}=37.35"